# Adaptive Federated Learning with Functional Encryption: A Comparison of Classical and Quantum-safe Options

### v1

@@ VI Experimental setup
In our experiments, we used the latest version of FLAD’s code at the time of writing (Release 1.0 [20]), executed in a Python 3.9 environment with Tensorflow version 2.7.1. All experiments have been performed on a server-class computer equipped with an AMD EPYC 9454P 48-Core Processor and 128 GB of RAM.
@@
Figure 2: Mean final accuracy and number of rounds relative to decimal digits truncation.
The experimental settings are consistent with those documented in the FLAD’s paper [2] in terms of dataset (CIC-DDoS2019 [21]), data preprocessing (an array-like representation of traffic flows, where the rows of the grids represent packets in chronological order and columns are packet-level features), unbalanced and heterogeneous data across the clients (one and only one attack assigned to each client) and ML model (a Multi-Layer Perceptron (MLP) with two hidden layers of $32$ neurons each, and an output layer with a single neuron for binary classification of the network traffic as either benign or DDoS). The total number of parameters of this model is 4 64146414\ 6414 641.
In order to achieve $128$ bits of security, we used Lattice Estimator [22] to learn the LWE parameters, along with the considerations in [11], and the NIST guidelines [23] for the DDH parameters.
Please note that all benchmarks and experiments are conducted using datasets available in the original FLAD repository [20]. We approximate floating points to integers by multiplying them for a common factor $10^{\Delta}$ and truncating the remaining digits. We set $\Delta=2$ since it is the most efficient approximation, achieving both small bitlength of the plaintexts and good performance in federated training, as in Figure 2. Data were collected on 11110 federated trainings with different random initial parameters, for each value of $\Delta$.
We also exploit parallelization to speed up both encryption and decryption of the models. We divided the plaintext (ciphertext, respectively) list into $15$ chunks with the same size. Then, we assigned a different process to each chunk.
@@
TABLE I: Parameters to achive $\approx{128}$ bits of security.
VII Experimental results
In this section we discuss the memory cost and the efficiency of the proposed scheme, by running a full training. We use a different cryptographic protocol in each run, to compare their performance. As we are about to see, despite the fact that LWE-based approaches grant post-quantum security, their selective version copes with DDH in terms of time consumption, but it is highly memory-consuming.
In Table II we compare the proposed schemes according to the memory cost. This comparison does not take into account the labels impact. Note that, since the vector y consists entirely of ones, each entry requires only one bit for storage.
$\mathbf{|cks_{i}|}$ $\mathbf{|sk_{y}|}$ $\mathbf{|ct|}$
@@ VII Experimental results
TABLE II: Memory cost DDH-based and LWE-based schemes.
Plugging the parameters shown in Table I in the formulas in Table II, a notable difference is the space required for each client’s private key in DDH-based schemes respect to LWE-based ones. In particular, the formers require approximately $1$ KiB, whereas LWE-based schemes need $3$ to $10$ MiB. What stands out more about this comparison is the space required for each ciphertext in the adaptive LWE-based scheme, which is approximately $285$ KiB, whereas in the other three schemes it is around $1$ KiB. In our test, with $13$ clients and 4 64146414\ 6414 641 parameters, we get that at each round the ciphertexts use around $16.5$ GiB of memory. This shows the poor scalability of this particular scheme compared to the others.
Figure 3 shows a comparison between the average time needed by the slowest client to train their model in a round of FLAD, the average time that the client uses to encrypt the model parameters and the aggregation time needed by the server to decrypt the parameters of the output model. It should be noted that in both DDH-based schemes, the aggregation algorithm requires performing a discrete logarithm. We solved it by precomputing a lookup table.
Figure 3 shows that the encryption in DDH-based schemes is more efficient compared to LWE-based schemes. When considering the total time required for a full round, the selective LWE-based scheme outperforms the DDH-based scheme, while the adaptive LWE-based scheme is notably less efficient. It is important to note that LWE-based schemes offer quantum resistance, a feature not shared by DDH-based schemes.
@@
VIII Conclusion and Future Work
In this paper, we have addressed both cyphertexts mix-and-mach and decryption-keys mix-and-match attacks in the context of Federated Learning (FL) applied to network security problems. To achieve this, we have integrated MIFE techniques with a recent adaptive FL implementation for network intrusion detection, namely FLAD.
While preventing reconstruction attacks, our experimental result show that adding a layer of cryptography heavily impacts the memory and time performances of FLAD. This is clear when the primitives are istantiated with security parameters that are appropriate for the modern standards. However, these findings give a realistic benchmark on the performances of two standard choices such as DDH-based and LWE-based MIFE, and represent a major improvement compared to similar studies in the scientific literature.
Surprisingly, the selective version of LWE-based MIFE can actually outperform its DDH counterpart, while preserving a moderate memory consumption and being post-quantum secure. Regarding the adaptive LWE-based MIFE, our work shows that it is not usable in a realistic scenario. This is fostering a deeper study on different quantum resistant solutions, such as RLWE (see [24] for more information).
In a further study, we want to compare the performances of different post-quantum secure primitives and to test and propose new batching routines that allows to process the client’s plaintexts at once. Another interesting aspect we aim to investigate is optimizations of MIFE that further exploit the specific features of FLAD.
Acknowledgment
This work was supported by Ministero delle Imprese e del Made in Italy (IPCEI Cloud DM 27 giugno 2022 – IPCEI-CL-0000007) and European Union (Next Generation EU).

### v2

@@ VII Experimental setup
In our experiments, we used FLAD Release 1.0 [26] in a Python 3.9 environment with Tensorflow version 2.7.1. All experiments were performed on a server with an AMD EPYC 9454P 48-Core Processor and 128 GB of RAM.
@@
Figure 1: Number of rounds and mean final accuracy over $10$ federated trainings with different random initial parameters for each decimal digit truncation $\Delta$.
The experimental settings are consistent with those documented in the FLAD’s paper [9] in terms of dataset (CIC-DDoS2019 [27]), data preprocessing (an array-like representation of traffic flows, where the rows of the grids represent packets in chronological order and columns are packet-level features), unbalanced and heterogeneous data across the clients (one and only one attack assigned to each client) and ML model (a Multi-Layer Perceptron (MLP) with two hidden layers of $32$ neurons each, and an output layer with a single neuron for binary classification of the network traffic as either benign or DDoS). The total number of parameters of this model is $4641$.
In order to achieve $128$ bits of security, we used Lattice Estimator [28] to obtain LWE parameters that meet the bounds in [12] Section 4.1, and NIST guidelines [29] for DDH parameters. The chosen parameters are shown in Table I.
To minimize computational cost we approximate floating points to integers by multiplying them by a common factor $10^{\Delta}$, with $\Delta=2$, and truncating the remaining digits. This choice is specific to our dataset. As shown in Figure 1, $\Delta=2$ is the most efficient approximation achieving both small bit length of plaintexts, leading to a faster MIFE, and good performance in federated training. We also exploit parallelization to speed up both the encryption and decryption of the models. We divided the list of plaintexts and ciphertexts into $15$ chunks of the same size, and assigned a different process to each chunk.
@@
TABLE I: Parameters to achieve $\approx{128}$ bits of security.
VIII Experimental results
We assess the memory and computational cost of the proposed scheme by running a full training. The primary goal of our tests is to evaluate the overhead introduced by the cryptographic layer on FLAD. Regardless of the cryptographic hardness assumption and the dataset, this overhead is linear in the number of training parameters. In fact, for both encryption and aggregation, each entry in the vector of model parameters undergoes the same operation individually. Consequently, conducting additional tests on different datasets would not provide significant new information.
In Table II, we compare the proposed schemes by memory cost. This comparison does not take into account the impact of round labels $\gamma_{t}$. Note that, since the vector y consists entirely of ones, each entry requires only one bit for storage.
$\mathbf{|csk_{i}|}$ $\mathbf{|sk_{y}|}$ $\mathbf{|ct|}$
@@ VIII Experimental results
TABLE II: Memory cost comparison.
Plugging the parameters shown in Table I in the formulas in Table II, a notable difference is the space required for each client’s private key in DDH-based schemes compared to LWE-based ones. In particular, the formers require approximately $1$ KiB, whereas LWE-based schemes need $3$ to $10$ MiB. What stands out more about this comparison is the space required for each ciphertext in the adaptive LWE-based scheme, which is approximately $285$ KiB, whereas in the other three schemes it is around $1$ KiB. In our test, with $13$ clients and 4 64146414\ 6414 641 parameters, we get that at each round the ciphertexts use around $16.5$ GiB of memory. This shows the poor scalability of this particular scheme compared to the others.
Figure 2 shows a comparison between the average time needed by the slowest client to train its model in a round of FLAD, the average time that the client uses to encrypt the model parameters and the aggregation time needed by the server to decrypt the parameters of the output model.
Figure 2 shows that the encryption in DDH-based schemes is more efficient compared to LWE-based schemes. When considering the total time required for a full round, the selective LWE-based scheme outperforms the DDH-based scheme, while the adaptive LWE-based scheme is notably less efficient. However, LWE-based schemes offer quantum resistance, a feature not shared by DDH-based schemes.
@@
IX Scheme generalizations
Our scheme can be adapted to FL schemes other than FLAD; the same principles can be applied to enhance the security of a broader class of FL algorithms, provided that the underlying scheme guarantees some properties required for compatibility with our cryptographic extension:
  1. P1
    in the aggregation step, the mean, either weighted or arithmetic, is performed on each client’s parameters;
  2. P2
    all clients send a model to the server at every round;
  3. P3
    the assignment of steps and epochs either does not depend on client secrets, but only on their aggregation; or it depends on secrets known only to the client running it, such as in LABEL:alg:_trainload.
FLAD satisfies (P1) and (P2). For other schemes, (P1) can be solved by adding a round of communication between each client and the server. In many FL schemes, e.g. FedAvg, the mean $w^{t}$ is performed on a different subset $C_{t}$ of the clients depending on the round $t$, that is
$w^{t}=\frac{\sum_{i\in C_{t}}k_{i}w^{t-1}_{i}}{\sum_{i\in C_{t}}k_{i}},$ (2)
where $k_{i}$ is the size of the $i$-th client’s dataset. This can be addressed in a naive way by changing the MIFE key in each round, exposing the scheme to decryption key mix-and-match attacks. Instead, we can ask for each client $i$ to share with the server the MIFE encryption of the quantity
$\delta_{i}=\begin{cases}&k_{i}\ \ \text{if $i$ trained in the current round,}\\ &0\ \ \text{otherwise.}\end{cases}$ (3)
Note that the following holds:
$\frac{\sum_{i\in C_{t}}k_{i}w^{t-1}_{i}}{\sum_{i\in C_{t}}k_{i}}=\frac{\sum_{i\in C}\delta_{i}w^{t-1}_{i}}{\sum_{i\in C}\delta_{i}},$ (4)
where $C$ is the set of all the clients. Thus, once the clients get the quantity $\sum_{i\in C}\delta_{i}$ through MIFE, they can send to the server the encryption of the quantity $\frac{\delta_{i}}{\sum_{i\in C}\delta_{i}}w^{t-1}_{i}$. Once the server decrypts this data with the usual decryption key associated with the vector $\mathbf{y}=(1,\dots,1)$, it finally gets the desired new model $w^{t}$.
Even if a FL scheme does not satisfy (P1, P2, P3), it is possible to slightly modify the scheme in such a way that
  • this modification does not impact performance, such as the number of rounds to achieve convergence or the accuracy of the final model, e.g. our modification in the epochs and steps assignment in FLAD (see LABEL:alg:_trainload);
  • the trade-off between security guarantees and performance loss satisfies the use-case.
X Conclusion and Future Work
In this paper, we have addressed security issues in the context of FL applied to network security problems. Specifically, we have integrated MIFE techniques with FLAD, a recent adaptive FL implementation for network intrusion detection. Moreover, we have discussed how our approach can be generalized to a broader class of FL schemes.
Despite preventing reconstruction attacks, our experimental results show that adding a layer of cryptography heavily impacts the memory and time performance of FLAD, when security parameters are set to modern standards. However, these findings give a realistic benchmark for the performance of two standard choices such as DDH-based and LWE-based MIFE, and represent a major improvement compared to similar studies in the scientific literature.
Surprisingly, the selective version of LWE-based MIFE can outperform its DDH counterpart, while preserving moderate memory consumption and being post-quantum secure. Our work shows that adaptive LWE-based MIFE is costly in a realistic scenario. This encourages further study on different quantum resistant solutions, such as RLWE [30].
In a further study, we want to compare the performance of different post-quantum primitives and to test and propose new batch encryption routines for client plaintexts.
We also aim to investigate an extension of this approach to Federated Learning algorithms in which only clients in a random subset send their models in each round.